OpenStudy (anonymous):

which represents the reference angle for 5pi/6

OpenStudy (neer2890):

pi=180 degree 5pi/6=5*180/6=5*30=150 degree

OpenStudy (anonymous):

the choices are a. 5pi/6 b. pi/3 c. pi/4 d. pi/6

OpenStudy (neer2890):

ok

OpenStudy (anonymous):

so is it a??

OpenStudy (neer2890):

no.. it is pi/6

OpenStudy (anonymous):

thanks and ur 100 % sure?

OpenStudy (neer2890):

check this out..: https://answers.yahoo.com/question/index?qid=20090308213216AANZGCx

OpenStudy (anonymous):

ow thanks a lot:)

OpenStudy (neer2890):

you can also check this: http://math.tutorcircle.com/trigonometry/how-to-find-reference-angles-in-radians.html#